OpenStudy (anonymous):

How do you solve 3 over x+2 equals 2 over x-2 using cross multiplication?

OpenStudy (anonymous):

Yes. The stuff it's dividing move it to the other side of the equal sign multiplying. 3*(x-2)=2*(x+2) Then multiply out 3x-6=2x+4 move the 2x left and the 6 to the right 3x-2x=4+6 x=10

OpenStudy (anonymous):

when you do cross multiplication this is what happens\[\frac{3}{x+2} = \frac {2}{x-2} \rightarrow 3(x-2) = 2(x+2)\]

OpenStudy (anonymous):

The answer is x=4 I got 10 though

OpenStudy (anonymous):

Nevermind. x=4 doesn't work

OpenStudy (anonymous):

I got ten above. And ten does work. 3*(10-2)=2*(10+2) 3*8=2*12 24=24

OpenStudy (anonymous):

2(x+2)=3(x-2) 3x-6=2x+4 3x-2x=4+6 x=10